OpenStudy (anonymous):

A+B = 5 pie/4, ( (1+cotA) ( 1 + cotB) )/(cotAcotB) = ?

hartnn (hartnn):

( (1+cotA) ( 1 + cotB) )/(cotAcotB) = (1+cot A + cot B + cot A cot B) / cotA cot B now separate the denominator.

OpenStudy (anonymous):

tanAtanB + TanA + tanB + 1

hartnn (hartnn):

what ?

hartnn (hartnn):

= (1+cot A + cot B)/cotAcot B + (cot A cot B) / cotA cot B = (1+cot A + cot B)/cotAcot B + 1

hartnn (hartnn):

that '1' will be a trouble...

hartnn (hartnn):

cot (A+B) =.... ?

hartnn (hartnn):

ok, lets work with tan tan(A+B) = tan 5pi/4 = ... ?

OpenStudy (anonymous):

tan 5pi/4=1

hartnn (hartnn):

tan(A+B) = (tan A+tan B) /(1+tan A. tan B) = ...

OpenStudy (anonymous):

tan(A+B) = (tan A+tan B) /(1-tan A. tan B)

hartnn (hartnn):

oops, tan(A+B) = (tan A+tan B) /(1-tan A. tan B) =1 which gives tan A+tan B = 1-tan A tan B

hartnn (hartnn):

so, tanAtanB + TanA + tanB + 1=...?

hartnn (hartnn):

tanAtanB + TanA + tanB + 1= tanAtanB +1-tan Atan B+ 1=2

OpenStudy (anonymous):

Yup... thxx...